A firm operating under perfectly competitive market has a cost function of TC= 1/3 Q3-Q2+10Q and market price =10 birr. Therefore, Find

1)     Level of output which maximizes profit of the firm

2)    Total Revenue(TR) 

3)     Total Cost(TC)

4)     Profit of the firm

5)    Shut down price and what is your decision as a manager based on your profit analysis.


1
Expert's answer
2021-01-19T07:50:33-0500

1) "Profit=P*Q-TC=10Q-(\frac{1}{3}*Q^3-Q^2+10Q)=Q^2-\frac{1}{3}*Q^3"

"MaxProfit=2Q-Q^2=0"

"Q1=0; Q2=2"

2) "TR=10*2=20"

3) "TC=19"

4)"Profit=20-19=1"

5) "19\/2=9.5" Shut down price

My decision is to produce 2 units